OpenStudy (anonymous):

f (x) = the square root of 7x + 42. I need to find the domain of the function. Could you help me please?

OpenStudy (anonymous):

\[f(x)=\sqrt{7x+42}\] requires that \(7x+42\geq 0\) can you solve that for \(x\)?

OpenStudy (anonymous):

subtract 42 to get \(7x\geq -42\) divide by 7 to get \(x\geq 6\)

OpenStudy (anonymous):

would it be x is greater than -6?

OpenStudy (anonymous):

thank you for helping me solve this equation